### v2

# FPGA-based Emulation and Device-Side Management for CXL-based Memory Tiering Systems

Abstract.
The Compute Express Link (CXL) technology facilitates the extension of CPU memory through byte-addressable SerDes links and cascaded switches, creating complex heterogeneous memory systems where CPU access to various endpoints differs in latency and bandwidth. Effective tiered memory management is essential for optimizing system performance in such systems. However, designing an effective memory tiering system for CXL-extended heterogeneous memory faces challenges: 1) Existing evaluation methods, such as NUMA-based emulation and full-system simulations like GEM5, are limited in assessing hardware-based tiered memory management solutions and handling real-world workloads at scale. 2) Previous memory tiering systems struggle to simultaneously achieve high resolution, low overhead, and high flexibility and compatibility.
In this study, we first introduce HeteroBox, a configurable emulation platform that leverages real CXL-enabled FPGAs to emulate the performance of various CXL memory architectures. HeteroBox allows one to configure a memory space with multiple regions, each exhibiting distinct CPU-access latency and bandwidth. HeteroBox helps assess the performance of both software-managed and hardware-managed memory tiering systems with high efficiency and fidelity. Based on HeteroBox, we further propose HeteroMem [^1] The code of HeteroBox and HeteroMem will be released after publication., a hardware-managed memory tiering system that operates on the device side. HeteroMem creates an abstraction layer between the CPU and device memory, effectively monitoring data usage and migrating data to faster memory tiers, thus hiding device-side heterogeneity from the CPU. Evaluations with real-world applications show that HeteroMem delivers high performance while keeping heterogeneous memory management fully transparent to the CPU, achieving a 5.1% to 16.2% performance improvement over existing memory tiering solutions.
* Co-first authors.$\dagger$ Corresponding author.
copyright: nonejournalyear: 2024doi: XXXXXXX.XXXXXXXconference: The 52nd IEEE/ACM International Symposium on Computer Architecture; June 21–25, 2025; Tokyo, Japanisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
The increasing demand for DRAM in modern data centers has led to its significant share in both operational costs and power consumption, consuming approximately 37% of the total cost and 33% of the power, according to Meta (Maruf et al., 2023). This dependency on DRAM is further entrenched by the DDR standard, which is designed around DRAM-specific access and maintenance commands, thereby limiting the integration of emerging memory technologies.
The emerging Compute Express Link (CXL) technique enables CPUs to access a variety of memory types by offering unified load-store semantics. This breakthrough paves the way for the adoption of alternative memory technologies, such as Managed DRAM (Lee et al., 2019), ReRAM (Chen, 2020), and 3D-XPoint/Optane (Hady et al., 2017), which promise larger capacities and lower costs (Dulloor et al., 2016) (Eisenman et al., 2018) (Kassa et al., 2021), potentially reducing the Total Cost of Ownership (TCO) for data centers (Sharma et al., 2024). As shown in Figure 1-(a), different types of memory media can be integrated into single or multiple CXL memory devices. CPU can directly access the expanded memory via CXL channels, avoiding modifying the DDR memory controllers in existing CPUs. In addition to enhancing the diversity of memory media connected to the CPU, the CXL standard supports connecting the host to multiple endpoint devices via switches, enabling larger-scale memory expansion. Additionally, CXL 3.0 introduces the capability for cascaded CXL switches (foundation, [n. d.]). As illustrated in Figure 1-(b), future CXL memory systems can feature a CXL switch connected to a host, another CXL switch, or a CXL memory device, thereby creating a complex topology.
The high flexibility of the CXL interconnect, while advantageous, also presents challenges in performance optimization. This duality is evident in several aspects. Firstly, new memory types (e.g., Optane) exhibit distinct latency and bandwidth characteristics. Integrating these memory media into a system via CXL results in a high heterogeneity. Secondly, the inclusion of CXL switches exacerbates such heterogeneity. As shown in Figure 1-(b), in a CXL-extended memory system with cascaded switches, different CXL memory devices may have varying numbers of CXL switches on the critical path to the host, resulting in different latency and bandwidth characteristics to the host CPU. In such systems featuring heterogeneous memory types and topologies, we should strive to ensure that the CPU’s memory access traffic takes the fast path rather than the slow one to preserve performance.
Figure 1. CXL-extended Heterogeneous Memory System
Prior research has endeavored to refine systems equipped with CXL memory from both system (Maruf et al., 2023; Lee et al., 2023) and architectural perspectives (Zhou et al., 2024; Zhong et al., 2024). The main idea of existing solutions is to leverage the access locality of workloads, identifying frequently accessed hot data and placing it into fast memory tiers. The remaining cold data is migrated to slow memory tiers to make room for hot data (Ren et al., 2023; Sha et al., 2023; Dong et al., 2010a). However, prior works face challenges in optimizing the performance of the emerging CXL-extended heterogeneous memory systems. The reasons are as follows:
Challenge 1: Lack of Evaluation Platform - Lacking real-world CXL-enabled CPUs and devices, previous studies rely on remote NUMA node emulation or software simulations to evaluate performance in CXL scenarios. However, NUMA emulation-based solutions cannot arbitrarily configure the latency and bandwidth characteristics of memory and are challenging to apply to hardware-managed memory tiering systems. Meanwhile, software simulation approaches struggle with handling large workloads and assessing server CPU performance at a real scale due to low simulation speeds. Several studies test commercial CXL devices, but the latency and bandwidth attributes of these devices are not configurable, and CXL switches are currently unavailable.
Challenge 2: Low Performance of Software-Managed Memory Tiering System - As detailed in Section 2, previous software-based memory tiering systems, which rely on various memory access profiling mechanisms including PTE scanning, hint-fault monitoring, and PMU sampling, struggle to accurately and efficiently identify hot and cold data due to hardware limitations. Meanwhile, the overhead of software-based data migration is significant.
Challenge 3: Poor Flexibility of Previous Hardware-Managed Memory Tiering System - Previous hardware-based memory tiering systems, primarily proposed before the CXL era, achieve memory access profiling and data movement by incorporating custom hardware logic in the CPU. These intrusive modifications to the CPU prevent these solutions from adapting to various CXL-extended memory system configurations.
Our Work: To address Challenge 1, we propose HeteroBox, an emulation platform based on a CXL-enabled FPGA. HeteroBox abstracts each CXL memory device as an addressable memory region that the host CPU can access with standard memory loads and stores. Each memory region can be assigned latency and bandwidth attributes corresponding to the emulated memory features. For Challenge 2, we propose HeteroMem, a hardware-managed memory tiering system that efficiently profiles data hotness and migrates data to faster memory tiers using hardware logic. HeteroMem includes a profiling unit that tracks every memory request from the host CPU and identifies hot and cold pages, forwarding them to a migration unit. The migration unit is optimized for efficient data movement with high internal bandwidth. For Challenge 3, we build HeteroMem logic on the device side, which acts as an abstraction layer between the CPU and device memory, hiding device-side heterogeneity from the CPU. A remapping unit in HeteroMem translates the address of memory request from host to device physical address, keeping migration completely transparent to CPU. Our main contributions are as follows:
  • We propose HeteroBox, a configurable emulation platform leveraging real CXL-enabled FPGAs to emulate the performance of various CXL memory architectures. By carefully designing its architecture, we enable HeteroBox to support multiple emulated regions, each with distinct latency and bandwidth characteristics (Section 3).
  • We propose HeteroMem, a hardware-managed memory tiering system operating on the device side. We carefully design the architecture of HeteroMem to accurately and efficiently detect hot and cold data, and to manage data placement through effective data migration (Section 4).
  • We evaluate HeteroMem’s performance on the HeteroBox emulation platform and explore various design choices for hardware-managed memory tiering systems. Our findings demonstrate that HeteroMem achieves a performance improvement of 5.7% $\sim$ 17.6% over several existing memory tiering systems, highlighting its effectiveness and potential as a superior memory management solution (Section 6).
2. Background & Motivation
Table 1. Comparison of HeteroMem with other Memory Tiering Systems
Memory Access Profiling Data Movement Flexibility and Compatibility
Mechanism Overhead Resolution Mechanism CPU Overhead Bandwidth CPU Transparent Runtime Configurable
AutoNUMA (Corbet, 2012) Hint-fault monitoring High Low Moved by CPU High Low No Yes
TPP (Maruf et al., 2023) Hint-fault monitoring High Low Moved by CPU High Low No Yes
DAMON (Park, 2021) PTE Scanning High Low Moved by CPU High Low No Yes
MEMTIS (Lee et al., 2023) PMU Sampling Median Median Moved by CPU High Low No Yes
HeMem (Raybuck et al., 2021) PMU Sampling Median Median Moved by CPU High Low No Yes
MemPod (Prodromou et al., 2017) Hardware Counters Low Median Moved by Hardware Low High Yes No
Neomem (Zhou et al., 2024) Hardware Hot Page Profiler Low High(only for hot pages) Moved by CPU High Low No Yes
HeteroMem Bidirectional Hardware Profiling Low High(both hot pages and cold pages) Moved by Hardware Low High Yes Yes
2.1. CXL-based Memory Expansion
The Compute Express Link (CXL) protocol(foundation, [n. d.]), based on the PCIe 5.0 physical layer, facilitates a cache-coherent, byte-addressable interconnect through high-efficiency SerDes links. This protocol composes three primary sub-protocols: CXL.io, CXL.cache, and CXL.mem. The CXL.mem sub-protocol is particularly noteworthy as it allows for direct CPU access to memory devices at cache-line granularity through the CXL interface. Unlike traditional DDR memory architectures with memory controllers in the host CPU, CXL incorporates memory controllers on the device side. This decouples the architecture, allowing flexible integration of various memory technologies into CXL memory controllers while preserving the host CPU design. Such flexibility is crucial for meeting specific capacity, performance, and cost-effectiveness needs of modern memory systems (Hynix, [n. d.]; Sumsung, [n. d.]; Li et al., 2023a; Boles et al., 2023).
CXL 2.0 introduces the CXL switch, which supports a single level of switching. The CXL switch allows one or multiple hosts to connect to one or more CXL devices. CXL 3.0 allows multiple switch levels, which enables more complex CXL memory device topologies. CXL 3.0 uses Port Based Routing (PBR) for inter-switch links to route CXL messages, allowing the software running on the host to perceive a flat topology ending at the leaf switch port. However, using a CXL switch introduces additional latency. According to (Li et al., 2023a), each CXL switch that a CXL request passes through adds about 70ns of latency. In a complex CXL memory system consisting of multiple switch levels, memory requests to different devices may pass through varying numbers of switches, resulting in a memory system with heterogeneous latency and bandwidth characteristics to the host CPU.
However, due to the limited availability of real-world CXL-enabled CPUs and devices, the evaluation of CXL-extended memory systems faces significant constraints. Specifically, previous work often relies on remote NUMA node emulation (Maruf et al., 2023; Lee et al., 2023) or software-based simulation methods (Jang et al., 2023) to assess the performance of CXL-based heterogeneous memory systems. But real CXL-extended memory exhibits different latency and bandwidth characteristics compared to remote NUMA nodes according to (Sun et al., 2023). Furthermore, CXL decouples the CPU and memory interface, allowing for more complex modifications on the device side, which cannot be modeled by NUMA emulation-based platforms. Software-based simulation methods, however, struggle with handling large workloads and assessing server CPU performance at a real scale due to low simulation speeds. For instance, the simulation speed of gem5 (Lowe-Power et al., 2020) in full system mode is approximately 100 kilo instructions per second (KIPS), which is tens of thousands of times slower than emulation methods. While some studies utilize CXL-enabled CPUs and devices (Sun et al., 2023; Zhang et al., 2023; Sano et al., 2023) for evaluation, their configurations are restricted by the available hardware and can hardly adapt to other settings. These limitations highlight the need for an emulation platform that can efficiently and accurately evaluate the performance of CXL-extended heterogeneous memory systems.
Insight#1: There’s an urgent need for a practical, high-fidelity method to evaluate the complex CXL-extended heterogeneous memory system.
2.2. Memory Tiering System
Many studies have focused on developing memory tiering systems to enhance the performance of heterogeneous memory systems. Specifically, we examine three critical aspects of memory tiering system design: (1) Memory Access Profiling, (2) Data Movement, and (3) Flexibility and Compatibility. Table 1 provides a summary comparison of prior work.
Figure 2. Overhead of Software Profiling and Data Movement
Memory Access Profiling: The memory access profiling mechanism is crucial for managing heterogeneous memory systems and achieving performance improvements. However, compared to hardware-based methods, software-based memory access profiling methods suffer from high overhead and low resolution. Software-based methods primarily consist of three types: hint-fault monitoring, PTE scanning, and PMU-based sampling. Hint-fault monitoring periodically poisons a portion of the memory pages, invalidating their PTEs and causing page faults when they are next accessed. TPP (Maruf et al., 2023) and AutoNUMA (Corbet, 2012) utilize hint-fault monitoring to detect hot data in slow memory. PTE scanning uses the access bit in the PTE to detect hot data. When a page is accessed, the MMU automatically sets the access bit. A kernel thread periodically scans the memory space to check whether a page has been accessed during the last period. AMP (Heo et al., 2022) uses PTE scanning to detect hot data in slow memory, while TMTS (Duraisamy et al., 2023) uses PTE scanning to detect cold data in fast memory. PMU sampling leverages specific hardware counters in the CPU to profile hot data in slow memory. For example, Intel CPUs support using PEBS counters to record LLC miss events. Each time an LLC miss event occurs, the PEBS counter increments by one. When the counter reaches a user-defined threshold, it overflows and records the access address. HeMem (Raybuck et al., 2021) and MEMTIS (Lee et al., 2023) use PMU sampling to detect hot data in slow memory.
However, as discussed in previous work (Zhou et al., 2024), these mechanisms fail to achieve both low overhead and high resolution simultaneously. As shown in Figure 2-(a), we use DAMON (Park, 2021) to profile the CPU overhead introduced by PTE scanning as an example. DAMON is a data access monitoring framework subsystem for the Linux kernel, which can be configured to profile memory access with different spatial and temporal granularities. We vary the space granularity from 50 regions to 1050 regions and the time granularity from scanning the page table every 50ms to every 1050ms. We visualize the overhead as a heat map, where lighter colors indicate heavier overhead. We observe significant overhead when using small spatial or temporal granularities, with a maximum cost of 88.15% CPU time.
Insight#2: Software-based memory access profiling methods suffer from high overhead and low resolution.
Figure 3. Overview of HeteroBox Emulation Platform.
Data Movement: After hot data is detected in slow memory, a data movement mechanism is required to migrate it to fast memory. Typically, the CPU can migrate data between fast and slow memory. Many previous works utilize the migrate_pages API in the Linux kernel for this purpose. However, as demonstrated in Nimble (Yan et al., 2019), using the CPU for data migration incurs significant overhead and reduced bandwidth due to the software stack in the OS, including tasks like checking permissions and unmapping and remapping pages. Although some works attempt to mitigate these limitations by using transactional or parallel migration, they still rely on the CPU for data migration, which inevitably causes CPU overhead. Furthermore, using the CPU to migrate data between CXL memory devices requires reading the data out and then writing it back, which consumes the limited bandwidth of the CXL link. As shown in Figure 2-(b), we compare the bandwidth of using custom hardware logic, as described in Section 4, to migrate data versus using the CPU to migrate data. For both methods, we use the 4KB page granularity and record the maximum number of pages that can be migrated per second. We observe that using custom hardware logic to migrate data achieves 12.9x higher bandwidth compared to using the CPU.
Insight#3: Using the CPU to move data results in higher overhead and lower bandwidth compared to device-side hardware-managed data movement.
Flexibility and Compatibility: Previous software-based CXL memory tiering systems can be classified into two categories. One approach is to present CXL memory to the CPU as NUMA nodes, where each node consists of homogeneous memory. The advantage of using NUMA abstraction is that it can seamlessly leverage previously optimized mechanisms in the operating system. Another approach is to use the DAX mode of CXL memory, in which the CXL memory device is exposed to the CPU as a file that supports load-store semantics. In this scenario, users must explicitly modify the source code of the application to use the DAX file as extended memory. Both using NUMA and DAX modes for heterogeneous memory management are not transparent to the CPU, reducing flexibility and compatibility.
Other works employ hardware-based memory management for better performance. For example, MemPod (Prodromou et al., 2017) modifies the CPU-side memory controller to group different types of memory into several pods. The MemPod logic in the memory controller optimizes the performance of the heterogeneous memory system without CPU involvement. These hardware-based heterogeneous memory management systems remain transparent to CPU. However, due to architectural limitations, they require modifications to CPU hardware, making them unsuitable to adapt to various configurations of CXL memory.
A recent state-of-the-art work, NeoMem (Zhou et al., 2024), proposes integrating a hardware unit within the CXL memory device, which communicates with the host CPU to provide memory access information. However, NeoMem relies on CPU to process this information and perform subsequent data movement, making it non-transparent to CPU and reducing overall flexibility and compatibility.
Insight#4: Previous software-based and hardware-based memory tiering systems fail to adapt to the complex configurations of CXL-extended memory systems.
To address these limitations, we first develop HeteroBox, an emulation platform designed to emulate a multi-tier CXL-extended heterogeneous memory system. Building on this platform, we introduce HeteroMem, a CXL-native, hardware-based memory tiering system. The core design philosophy of HeteroMem is to create an abstraction layer between host CPU and memory controller within the CXL device. This hardware-based abstraction layer is tasked with identifying hot data in slower memory and migrating it to faster memory. This entire process is completely transparent to host CPU. Further details are introduced in the following sections.
3. HeteroBox Platform
3.1. Overview of HeteroBox Emulation Platform
Figure 3 illustrates the overview of HeteroBox emulation platform, where a CXL-enabled host processor (e.g., Intel 4th-generation Xeon Sapphire Rapids CPU) connects to a CXL FPGA board that manages the device memory. We implement the emulation logic of HeteroBox between the memory controller and CXL controller, intercepting and processing all memory read requests. In the following description, we refer to the request address in the host address space as hPA and the address sent to the memory controller in the device as dPA. Typically, hPA equals dPA when no additional translation logic is involved. The requests are assigned different latency and bandwidth characteristics according to their dPA. Besides, we implement configuration registers for HeteroBox, mapping them to the PCIe BAR space to control the behavior of HeteroBox. A host-side software interface allows users to configure HeteroBox at runtime. Meanwhile, users can implement custom hardware logic (e.g., hardware-based tiered memory management unit) on the HeteroBox emulation platform.
Using a CXL-enabled CPU and FPGA, HeteroBox accurately emulates the performance characteristics of the CXL link, and its emulation logic replicates various latency and bandwidth characteristics. These capabilities allow HeteroBox to precisely model a wide range of CXL-extended heterogeneous memory systems. In the subsequent part of this section, we will detail the design of HeteroBox’s emulation logic and its configuration process.
3.2. Emulation Logic Design
The emulation logic of HeteroBox creates an illusion that the homogeneous DDR memory on the FPGA is divided into several heterogeneous parts, each with distinct latency and bandwidth characteristics. As shown in the right part of Figure 3 (), the emulation logic contains a timestamp register, and for each abstract region, there is a latency register, a bandwidth configuration register and a bandwidth counter register. The timestamp register increments by one each cycle, while the latency registers and bandwidth configuration registers can be configured by users through CXL.io requests. We focus on read requests since they are on the critical path of program execution. When the ready signal is high, a read request can enter the emulation logic. HeteroBox first identifies the memory region of the dPA for the read request. It then attaches a timestamp, incremented by the latency register of the corresponding memory region, to the request as a tag. In practice, HeteroBox does not send the tag to the memory controller; instead, it pushes the tag into a FIFO and maintains the correspondence between each tag and the request’s response. When the read response is returned from memory, it is first pushed into a response FIFO. HeteroBox then checks the tag of the first element in response FIFO every cycle to determine if the value of the timestamp register exceeds the tag. If it does, HeteroBox pops the response out and returns it, adding a configurable latency to the region.
For bandwidth characteristics, HeteroBox maintains a bandwidth configuration register and a bandwidth counter register for each region. The bandwidth configuration register is user-configurable, while the bandwidth counter register increments by one whenever a response is returned from the corresponding region, and it resets at a user-defined interval. When the bandwidth counter register reaches the value set in the bandwidth configuration register for that region, HeteroBox blocks further responses from that region until the bandwidth counter register is reset in the next interval.
3.3. Configuration Process
As shown in the left part of Figure 3 (), the configuration registers for HeteroBox are mapped to the PCIe BAR space of the host. These registers control the number of memory regions, the dPA range of each region, and their latency and bandwidth characteristics. The host CPU can read or write these registers through MMIO.
The HeteroBox platform supports multiple regions with distinct latency and bandwidth characteristics. To configure HeteroBox, we implement drivers in kernel space and an interpreter in user space. Users can configure HeteroBox using JSON files, as illustrated in the left part of Figure 3 (). These files specify the number of memory regions, region IDs, start and end dPAs, latency (cycles at 200MHz), and bandwidth (bandwidth configuration register). The JSON file is processed by user space interpreter, which sends the information to the kernel space driver. The driver then writes to PCIe BAR space to update configuration registers in HeteroBox.
4. HeteroMem Solution
4.1. Overview of HeteroMem System
Figure 4. System Overview of HeteroMem
The HeteroMem memory tiering system acts as an intermediate layer between host CPU and CXL-extended memory. As shown in Figure 4, the HeteroMem system comprises three main components: the Remapping Unit, the Profiling Unit, and the Migration Unit. Additionally, we provide software interface support for HeteroMem memory tiering system, including a software driver in the kernel and BAR registers in the device for configuration and profiling. For the following evaluation of HeteroMem memory tiering system, we use a two-tier memory system, consisting of a fast memory tier and a slow memory tier, without losing generality.
Figure 5. Remapping Unit Working Process.
Figure 6. Block Diagram of the Remapping Unit
Remapping Unit: HeteroMem manages heterogeneous memory system entirely on the device side, migrating hot data to fast memory and cold data to slow memory. To keep this migration transparent to CPU, the device needs a translation layer to map hPA to dPA. We implement a Remapping Unit to handle this, translating memory requests and forwarding them to subsequent modules. Since the Remapping Unit is on the critical path of memory requests, it must minimize added latency to ensure optimal system performance.
Profiling Unit: The Profiling Unit assesses the hotness and coldness of data in CXL-extended memory using translated memory requests from the Remapping Unit. It records access patterns, profiling hot data in slow memory and cold data in fast memory, temporarily buffering the dPA of detected cold data. When hot data is identified, the Profiling Unit signals Remapping Unit to initiate a migration event, providing the dPA for hot data and the dPA of previously buffered cold data for subsequent migration transaction.
Migration Unit: The Migration Unit initiates data migration upon receiving requests from the Remapping Unit, which contain dPAs for hot and cold data identified by the Profiling Unit, and swaps their memory locations. To keep migrations transparent to the CPU, the Remapping Unit must be updated whenever a migration occurs.
4.2. Remapping Unit
The Remapping Unit provides an additional translation layer between the CPU and the device. It needs to be carefully designed to introduce minimal overhead and keep the translation process transparent to the CPU.
Memory Layout: As shown in Figure 5-(a), the entire memory space consists of a fast memory region and a slow memory region. The Remapping Unit stores the metadata for address translation at the beginning of the fast memory, which is software-reserved to avoid access by the host CPU. When the machine powers on, the Remapping Unit initializes this metadata by sending a series of write requests to the reserved fast memory. This metadata includes a remapping table array and a reverse remapping table array, both consisting of 4-byte entries, each containing a page index (for dPA). HeteroMem manages memory in 4KB page granularity, so a 4-byte page index can describe a memory space of up to 16TB. For larger memory spaces, a longer index (e.g., 8 bytes) can be used. Accessing the remapping table array with a page index yields the translated page index, while the reverse remapping table contains the inverse mapping of the remapping table.
Architecture of Remapping Unit: The Remapping Unit’s architecture, shown in Figure 6, works as follows: when a memory request arrives, it first checks the remapping cache, which stores metadata for translating requests. If the required metadata is found, the request is translated and sent to memory. On a cache miss, the request is buffered in a FIFO, and the remapping cache issues a memory read to retrieve the needed remapping table data. Once the data returns, the cache is updated, and the buffered request is translated and forwarded to memory. Figure 5-(b) illustrates the translation process of the Remapping Unit when a remapping cache miss occurs.
Figure 7. Profiling Unit Overview.
Migration Transaction: As shown in Figure 5-(c), when the Remapping Unit receives a migration request from the Profiling Unit (❶), it initiates a migration transaction. The migration request contains a pair of page indexes located in fast memory and slow memory, respectively. The page indexes in the migration request have already been translated. Upon receiving the migration request, the Remapping Unit blocks subsequent requests from the host to ensure the atomicity of the migration. Then, the Remapping Unit issues two memory read requests to the reverse remapping table in the memory to retrieve the hPAs of the hot and cold page (❷). When the read responses containing the hPAs return (❸), the Remapping Unit updates the remapping table data and the reverse remapping table data according to the two hPAs and two dPAs (❹). The Remapping Unit concurrently sends an enable signal to the Migration Unit after issuing the memory read request to overlap the memory read latency. The Migration Unit then reads the data of the two pages at the two page indexes specified by the migration request (❻❼), swaps their indexes, and writes them back (❽❾). After the Migration Unit has completed the migration process and the Remapping Unit has updated its remapping and reverse remapping table data, subsequent memory requests are allowed to proceed.
4.3. Profiling Unit
The Profiling Unit is responsible for profiling the hotness and coldness of data in memory. In our HeteroMem design, we profile data hotness and coldness at a 4kB page granularity. The Profiling Unit is linked after the Remapping Unit, off the critical path, to capture the dPA trace. We profile only read requests, as write requests are generally not on the critical path of program execution. As shown in Figure 7-(a), in the case of a CXL-extended memory system with two tiers, the Profiling Unit splits the read requests into two streams: one targeting fast memory and the other targeting slow memory. For the read request stream targeting fast memory, the Profiling Unit detects cold pages based on their access history. These detected cold pages are temporarily buffered in a cold pages buffer. Since cold pages tend to remain cold for a relatively long time (Duraisamy et al., 2023), they typically remain cold when fetched from the buffer in subsequent processes. For the read request stream targeting slow memory, the Profiling Unit measures the hotness of each page based on its access frequency, reporting a page as hot if its hotness exceeds a given threshold. Upon detecting a hot page, the Profiling Unit fetches a cold page from the cold pages buffer and sends the detected hot page and the fetched cold page to the Remapping Unit. The Remapping Unit then manages the data movement, swapping pages between fast memory and slow memory. The number of these hot and cold page pairs sent to the Remapping Unit is limited by a given constraint (e.g., 32 pairs of pages per 100,000 cycles).
[…]
Coldness Profiling: As detailed in Section 4.D, the Migration Unit uses a swap mechanism to move hot data to fast memory. Therefore, when a hot page is detected in slow memory, we need to find a cold page in fast memory to swap its location with the detected hot page. We use a ping-pong bitmap to record the access history of each page in fast memory. The ping-pong bitmap structure consists of two bitmap arrays, each entry being a bit corresponding to a page in fast memory. In a given period, one bitmap, say bitmap $A$, is used to record memory access in the current period. When Profiling Unit receives a memory request, it sets the bit corresponding to the accessed page in bitmap $A$. Meanwhile, Profiling Unit scans the other bitmap, say bitmap $B$, to identify pages with their corresponding bit in bitmap $B$ unset as cold pages. In the next period, we reset all bits in bitmap $B$ and switch the functions of bitmap $A$ and bitmap $B$. This approach ensures the detected cold pages were not accessed during the last period, thus accurately identifying cold pages.
4.4. Migration Unit
The Migration Unit moves data between fast memory and slow memory. On the device side, we lack knowledge about whether an address contains valid data, so every migration operation is executed as a swap operation, swapping the data between fast memory and slow memory. The Migration Unit takes in a pair of dPAs and an enable signal. When the enable signal is set high, the Migration Unit stages the two input dPAs, reads the data from these dPAs into a buffer, and then writes the data to the swapped dPAs to complete the process. The Migration Unit issues read requests to the memory controller in a non-blocking manner and initiates write requests as soon as a read response returns, while simultaneously blocking incoming memory requests from the host during migration, thereby minimizing idle cycles and maximizing migration bandwidth.
4.5. Software Interface
HeteroMem can manage the memory tiering system transparently without CPU intervention. To better understand the behavior of HeteroMem, we have implemented a series of profiling registers, each indicating the number of specific events. Additionally, HeteroMem’s behavior can be tuned through several parameters, which we have also implemented as configuration registers. We expose these configuration and profiling registers in the PCIe BAR space of the CXL memory device. The host CPU can read or write these registers through MMIO. In the OS kernel space, we have implemented a device driver that exposes the MMIO address of these registers to user space as files under the $/sys/kernel/mm/heteromem$ directory. Users can read or write these files to configure HeteroMem and obtain profiling information. For example, users can read $/sys/kernel/mm/heteromem/migrated\_pages\_cnt$ to get the number of pages migrated by HeteroMem.
5. Implementation
Table 2. Evaluation System Configuration
Host CPU Single socket Intel® Xeon 6430 CPU @ 2.10GHz
32 Cores, hyperthreading disabled.
60MB Shared LLC
DDR Memory 32GB DDR5 4800MHz x 2
[…]
Hard CXL 1.1 IP on PCIe Gen5 x16
16GB 2-Channel DDR4-2666 DRAM
Software Environment Linux kernel version v6.3
Table 3. Parameters Configuration
Parameter Value Description
Sketch Width(W) 512k Width of the sketch in hotness profiling logic.
Sketch Depth(D) 2 Depth of the sketch in hotness profiling logic.
sketch_reset_period $50$-$5000ms$ The reset period of the sketch in HeteroMem.
bitmap_reset_period $1$-$5s$ The reset period of the bitmap in HeteroMem.
migrate_page_limit $256MB/s$ The limit of migration speed for all memory tiering systems.
pebs_sampling_rate $397$-$19997$ The sampling rate of PEBS
pte_sampling_rate $1$-$3s$ PTE-sampling rate of TPP&AutoNUMA
page_scanning_rate $5$-$10s$ Page table scanning rate of PTE-scan
Figure 8. FPGA-based Prototyping System
[…]
We implement HeteroMem based on the Linux kernel v6.3. The CXL memory on the FPGA is exposed to the software as a CPU-less NUMA node. To compare the performance of HeteroMem with other baseline memory tiering systems, we modified the NUMA nodes enumeration code and split the CXL memory NUMA node into two fake NUMA nodes, corresponding to the fast memory and slow memory emulated by HeteroBox. We use Intel® FPGA CXL IP (Memory Expander Type 3) and add about 7000 lines of Verilog RTL custom logic code to build HeteroBox and HeteroMem, along with approximately 1000 lines of C++ simulation environment code and about 1000 lines of modifications in the Linux kernel for HeteroBox and HeteroMem driver and fake NUMA nodes logic. We will open-source all the hardware and software codes after publication.
Figure 9. Validation of HeteroBox Emulation Platform
Our FPGA prototype of HeteroMem consumes 224.4k ALMs (24.6%) and 3615 BRAMs (M20K, 27.2%), with no usage of DSPs. The layout of our HeteroMem FPGA prototype is illustrated in Figure 8-(a), where the red part represents our HeteroMem logic. This logic acts as an abstraction layer between Intel CXL hard IP and memory controller. Figure 8-(b) provides a photo of our implementation platform. Finally, Figure 8-(c) presents an overview of the system emulated by HeteroBox for the subsequent evaluation. We configure HeteroBox to emulate two memory regions in CXL-extended memory space: one fast memory region and one slow memory region. Application memory is allocated in CXL-extended memory space, while OS memory remains in the host memory.
6. Evaluation
6.1. Experimental Setup
Baselines: We compare the performance of our HeteroMem design with five baseline memory tiering systems: PTE-Scan (Park, 2021), PEBS (peb, [n. d.]), Auto NUMA (Corbet, 2012), TPP (Maruf et al., 2023), and NeoMem (Zhou et al., 2024). PTE-Scan utilizes the access bit in the page table entry for hotness profiling. The access bit is set by the MMU when a page is accessed, and PTE-Scan periodically scans the page table, clearing the access bit and detecting a page as hot if its access bit is set. PTE-Scan migrates hot pages to fast memory as soon as they are detected. PEBS uses hardware counters to profile page hotness. It is configured to sample the $MEM\_LOAD\_L3\_MISS\_RETIRED$ event. When an LLC miss occurs, the hardware counters increment by one. When the counter reaches a given threshold, it overflows and records the address of the current memory access. The PEBS memory tiering system detects a page as hot and migrates it to fast memory when its address is recorded. TPP and AutoNUMA utilize hint-fault monitoring to detect hot data in slow memory. NeoMem is a recently proposed memory tiering system that achieves state-of-the-art performance. It enhances the CXL memory device with a hot page profiling structure, which identifies and reports hot pages to the host CPU for migration to faster memory. Table 3 presents the parameter configurations for these memory tiering systems.
Benchmarks: We select eight benchmarks for evaluation: GUPS (gup, 2021), a microbenchmark that performs parallel read-modify-write operations in a uniform or skewed random pattern within its working set; Silo (Tu et al., 2013), an in-memory database engine; Btree, an in-memory index lookup; XSBench (Tramm et al., 2014), an HPC workload; Betweenness Centrality (BC) and PageRank (PR)(Beamer et al., 2017), two graph processing workloads; and 603.bwaves and 654.roms, two benchmarks from SPEC 2017. These benchmarks are widely used to evaluate memory tiering systems in previous works(Raybuck et al., 2021; Lee et al., 2023; Zhou et al., 2024). The Resident Set Size (RSS) of these benchmarks ranges from 4.4GB to 11.1GB.
6.2. Validation of HeteroBox
We use the Intel MLC tool (mlc, [n. d.]) to evaluate the latency and bandwidth attribute of the HeteroBox emulation platform. We first focus on the case of a single region. As illustrated in Figure 9-(a), we configure the HeteroBox to contain one region and set the latency register value from 0 to 256. Using the Intel MLC, we then measure the region’s latency. Changing the bandwidth configuration values from 128 to 32, we plot four curves of relationship between the value of latency register and the latency attribute of this region. Additionally, we include a curve called “theoretical latency”, which incorporates the latency from both the CXL link and the CXL IP on the FPGA. The theoretical latency also includes the latency introduced by the HeteroBox emulation logic. Due to the latency introduced by the underlying memory controller and DDR4 DRAM, the region’s latency is the same when the latency register is set to 0 or 32. All four curves corresponding to different bandwidth configuration values overlap, showing consistent latency performance when compared to theoretical latency. These results demonstrate two key findings: (1) The HeteroBox emulation logic can precisely adjust latency attribute of a region. (2) Modifying the bandwidth configuration register value does not affect the latency attribute of the region.
As shown in Figure 9-(b), we plot five curves representing the relationship between the bandwidth configuration register value and the bandwidth attribute of the region. Each curve corresponds to a different latency register value. Additionally, a theoretical bandwidth curve is drawn by using the point where the bandwidth configuration register value is 32 as a base point, and expanding proportionally with the bandwidth configuration register value. The results demonstrate that: The HeteroBox emulation logic can accurately adjust the bandwidth attribute of a region.
Figure 10. Analysis of Remapping Overhead.
Figure 11. Performance Comparison with Different Fast Memory Size Configuration.
Next, we discuss the case of multiple regions. As shown in Figure 9-(c) and Figure 9-(d), we configure the HeteroBox with four regions, dividing the FPGA memory into four equal parts. We modify the kernel to create four NUMA nodes corresponding to these regions and use Intel MLC to evaluate their latency and bandwidth attributes. In Figure 9-(c), the bandwidth configuration register value for all regions is set to 128, while the latency register values for regions 1, 2, and 3 are set to 32, 64, and 96, respectively. We vary the latency register value of region 4 and plot the latency attributes of all four regions. In Figure 9-(d), the latency register value for all regions is set to 128, while the bandwidth configuration register values for regions 1, 2, and 3 are set to 128, 80, and 32, respectively. We vary the bandwidth configuration register value of region 4 and plot the bandwidth attributes of all four regions. The results show that: (1) The HeteroBox emulation logic can accurately adjust the latency and bandwidth attributes in the case of multiple regions. (2) Adjusting the latency or bandwidth of one region does not affect the attributes of the other regions.
6.3. Remapping Overhead
Then we discuss the overhead introduced by HeteroMem. The Profiling Unit of HeteroMem is off the critical path, so it does not introduce extra latency. The Migration Unit only operates when a migration transaction occurs. Since we limit the number of pages migrated within a given period (cycles consumed on migration are limited to below 5%), the extra latency introduced by Migration Unit is minimal. However, Remapping Unit is always on the critical path and incurs non-negligible extra latency and bandwidth overhead when the remapping cache misses. As a result, the overhead introduced by the Remapping Unit should be carefully considered.
As shown in Figure 10-(a), we disable migration in HeteroMem and test the performance slowdown with the Remapping Unit enabled, varying the remapping cache size from 128KB to 2MB. Figure 10-(b) displays the remapping cache hit rate for different cache sizes. For all benchmarks, both the performance and the remapping cache hit rate improve as the remapping cache size increases. The GUPS benchmark shows a low remapping cache hit rate with smaller cache sizes, causing significant overhead due to its random memory access pattern. This results in a low cache hit rate when the cache cannot cover the accessed memory range. However, as the remapping cache size increases and covers the memory range accessed by GUPS, the hit rate improves significantly, reaching nearly 100% at 2MB, reducing the Remapping Unit’s overhead. For BC, PR and XSBench benchmarks, the remapping cache hit rate also approaches 100%, resulting in minimal overhead with a 2MB cache. Conversely, the Btree and Silo benchmarks exhibit lower locality, leading to low remapping cache hit rates even with a 2MB cache, and show performance slowdowns of 11.8% and 11.0%, respectively. Results from subsequent experiments indicate that the performance degradation in these two applications caused by Remapping Unit can be offset by HeteroMem’s superior memory tiering solution.
6.4. Main Results
Performance Comparison: To evaluate the performance of HeteroMem, we configure a 2MB remapping cache and compare HeteroMem’s end-to-end performance with baseline systems. We run HeteroMem on HeteroBox emulation platform, dividing the memory space into fast and slow regions. The fast memory region size varies from 4GB to 1GB with no additional latency. The slow memory region has an additional latency of 128 cycles in a 200MHz clock domain. Neither memory region has any bandwidth restrictions.
Figure 12. Sensitivity Analysis and Profiling Result of HeteroMem
Figure 13. Analysis of Memory Access Distribution.
As shown in Figure 11, HeteroMem consistently maintains superior geomean performance compared to baseline systems across various fast memory sizes. All performance statistics are normalized to NeoMem. In the 4GB fast memory configuration, HeteroMem’s geomean performance is 5.1% to 16.2% higher than baseline systems. For BC benchmark, HeteroMem achieves 32.7% higher performance than NeoMem and PTE-Scan, and 44.6% higher than PEBS. Additionally, the performance improvement of HeteroMem increases with larger fast memory sizes, primarily due to the enhanced effectiveness of HeteroMem’s accurate cold page profiling mechanism. The performance improvement of HeteroMem stems from its accurate hardware-based profiling method and efficient hardware-based data movement mechanism, as described in Section 4.
For the following experiments, we utilize a configuration with 4GB fast memory and a 2MB remapping cache to delve into the superior performance of HeteroMem.
Different Slow Memory Latency: To explore how slow memory latency affects the performance of HeteroMem, we vary the additional latency in slow memory from 128 cycles to 256 cycles (200MHz clock domain) and test the performance of BC benchmark. The results are shown in Figure 12-(a). HeteroMem maintains superior performance compared to other memory tiering systems, with non-significant performance degradation as slow memory latency increases. This is because HeteroMem accurately and timely migrates hot pages from slow to fast memory. PEBS performs the worst because it samples hot pages, leading to fewer detected hot pages and less page migration. TPP also shows low performance due to its hysteretic promotion mechanism.
Number of Migrated Pages: As shown in Figure 12-(b), we profile the number of migrated pages for HeteroMem and five baseline systems. For the BC benchmark, all hot pages can fit in fast memory, so HeteroMem stops migrating once they are placed there. NeoMem, PTE-Scan and Auto NUMA show redundant migrations, while PEBS and TPP fail to migrate all hot pages due to slow convergence. In other benchmarks, where hot pages cannot all fit in fast memory, HeteroMem migrates more pages than the baseline systems. With a 256MB/s migration limit and higher bandwidth, HeteroMem’s additional migrations have negligible overhead. This results in a higher proportion of fast memory accesses, leading to improved end-to-end performance of HeteroMem.
Memory Access Distribution: As shown in Figure 13-(a), we profile the number of memory accesses to fast and slow memory for all six memory tiering systems. We show the total number of memory accesses and the ratio of memory accesses to slow memory. For all benchmarks except GUPS, HeteroMem shows fewer total memory accesses compared to the baseline systems due to the additional CPU memory accesses required by software-based data migration in the baseline systems. In the GUPS benchmark, HeteroMem exhibits more memory accesses because the performance metric of GUPS is based on the memory access count within a given period. Additionally, HeteroMem demonstrates a lower ratio of slow memory accesses across all benchmarks (except for GUPS) compared to the baseline systems. The higher proportion of fast memory accesses explains HeteroMem’s superior performance, showing its effective profiling of hot and cold data and optimal placement of hot data in fast memory.
In Figure 13-(b), we plot the memory access distribution for the BC benchmark. We implement a 32-bit hardware counter for each 2MB physical page in the CXL extended memory and record the memory access count to each 2MB page every second. The results show that HeteroMem quickly migrates all hot data to fast memory, whereas the baseline systems leave some hot data in slow memory, demonstrating the effectiveness of HeteroMem’s memory management strategy.
7. Related Work
7.1. Software-based memory tiering system
Many previous works using software-based methods have tried to optimize the performance of memory tiering systems in various aspects, including memory access profiling (Agarwal and Wenisch, 2017; Choi et al., 2021; Lee and Eom, 2022; Park, 2021), page migration (Kommareddy et al., 2019; Ryoo et al., 2018; Yan et al., 2019), page classification (Koh et al., 2019; Maruf et al., 2022), and huge page strategies (Heo et al., 2022; Lee et al., 2023). TPP (Maruf et al., 2023) proposes to promote a page after it is scanned with the access bit set twice. MEMTIS (Lee et al., 2023) uses histogram-based hot set classification to achieve more stable hot page detection. NOMAD (Xiang et al., 2024) employs transactional page migration to mitigate the overhead of page migration. Additionally, there are works that utilize application-level information to manage data placement at the object level (mem, 2021; pmd, 2017; Hildebrand et al., 2020; Li and Wu, 2022; Ren et al., 2021; Wang et al., 2019; Wei et al., 2015). However, although these works strive to optimize the performance of memory tiering systems, their performance is limited by the inherent inaccuracy of software profiling and the high overhead of data migration using the CPU, as described in Section 2. Our HeteroMem design employs a hardware-based high-resolution profiling scheme for hot and cold page detection. Meanwhile, HeteroMem uses hardware-based data movement, which achieves high bandwidth and low overhead without wasting CPU cycles and CXL link bandwidth.
7.2. Hardware-based memory tiering system
There are also many works that try to optimize memory tiering systems using hardware-based methods. They manage fast memory as a memory cache (opt, [n. d.]) or use fast memory in flat memory mode (Chou et al., 2014; Ryoo et al., 2017; Kotra et al., 2018; Dong et al., 2010b; Prodromou et al., 2017; Li et al., 2023b). Intel Optane persistent memory provides memory mode (opt, [n. d.]), allowing users to use DRAM as a cache for non-volatile memory. CAMEO (Chou et al., 2014) combines memory cache mode and flat address mode. It manages fast memory in a cache-like manner, constrains data movement within associated sets, and moves data at cacheline granularity. MemPod (Prodromou et al., 2017) groups HBM and DRAM connected to CPU into clusters called Pods and manages data placement within each Pod. However, these hardware-based methods, proposed before the CXL era, involve intrusive modifications to CPU, making them inflexible and unable to adapt to dynamic and diverse CXL-based device-side heterogeneous memory systems. HeteroMem, on the other hand, places all its hardware logic on the device side and can be runtime configured to adapt to various heterogeneous memory system configurations, offering a one-for-all solution.
A recent work, NeoMem, introduces a method to profile hot data on the device side, achieving state-of-the-art performance. However, NeoMem relies on CPU-side profiling for cold data and uses CPU to manage data movement, while HeteroMem employs device-side hardware logic for profiling both hot and cold data, as well as managing data movement. The design of HeteroMem creates a homogeneous abstraction for the CPU and improves overall performance compared to NeoMem.
8. Discussion & Future Work
Support for CXL memory pooling and sharing: Memory pooling and sharing are promising use cases of CXL (Ha et al., 2023), allowing memory to be shared between processors (consortium, [n. d.]). Memory in a memory pool can be dynamically allocated and deallocated, leading to more efficient use of memory resources. Memory sharing, on the other hand, enables each processor to read from or write to the shared memory, facilitating easy data sharing and communication between processors. The data migration process of HeteroMem is completely transparent to CPU. As a result, HeteroMem serves as a plug-in solution for both memory pooling and sharing scenarios.
Hardware overhead and Scalability: The hardware overhead of HeteroMem remains stable as the CXL-extended memory system scales. Firstly, HeteroMem’s hardware overhead increases minimally with the number of CXL devices. HeteroMem can be integrated into a CXL switch, and the number of CXL switches is significantly smaller than the number of CXL devices. Furthermore, not every CXL switch needs to be equipped with HeteroMem hardware. HeteroMem can manage the entire memory space it can access, so it only needs to be implemented in the CXL switch connected to the host CPU. Secondly, as the memory space scales, the increase in HeteroMem’s hardware overhead remains minimal. The hot page profiling component of HeteroMem utilizes Count-Min Sketch to detect hot pages, with an error bound related to the memory access sequence length rather than the memory space size. The cold page profiling component, namely the ping-pong bitmap, uses two bits of SRAM for every page. This structure only needs to be allocated for fast memory regions with demotion targets, which typically have a small capacity. For the Remapping Unit, the page granularity can be increased as memory space expands, such as using a mix of 4KB and 2MB pages, or exclusively 2MB pages. For example, a 4MB Remapping Cache (potentially larger in ASIC designs) with 2MB page granularity has a reach of 1TB memory (assuming 8 bytes per entry). This reach can be further extended by using larger page granularity (e.g., 1GB pages).
Profiling and Migration Granularity: In our current HeteroMem design, we use a 4KB page for both profiling and migration granularity for design simplicity. Generally speaking, smaller granularity enables more flexible data placement but introduces heavier remapping overhead. Using larger granularity reduces the overhead of remapping but introduces amplification in data movement and wastage of fast memory due to hotness skew in a large page. There are no theoretical difficulties in implementing profiling and migration granularities other than 4KB or using multiple granularities in HeteroMem. We leave this for our future work.
9. Conclusion
In this paper, we introduce HeteroBox, an innovative emulation platform built on a CXL-enabled FPGA. HeteroBox supports emulating a CXL-extended heterogeneous memory system with multiple regions, each with configurable latency and bandwidth characteristics. Based on the HeteroBox emulation platform, we propose HeteroMem, a novel hardware-based, device-side memory tiering system. HeteroMem is transparent to the host CPU and can be runtime configured to adapt dynamically to different CXL-extended heterogeneous memory system configurations. Additionally, HeteroMem achieves high performance through accurate memory access profiling and efficient data movement. Our evaluation results show that HeteroMem achieves a 5.1% to 16.2% performance improvement over existing memory tiering solutions.